How do I reduce w^9/w^12?
Answer by dabanfield(803)

How do I reduce w^9/w^12?
Since 1/w^12 = w^(-12):
w^9/w^12 =
w^9*w^(-12) = (remember (x^a)*(x^b) = x^(a+b))
w^(9-12) =
w^(-3) =
1/w^3